# Large Language Models for Computer-Aided Design: A Survey

### v1

@@ Abstract.
Large Language Models (LLMs) have seen rapid advancements in recent years, with models like ChatGPT and DeepSeek, showcasing their remarkable capabilities across diverse domains. While substantial research has been conducted on LLMs in various fields, a comprehensive review focusing on their integration with Computer-Aided Design (CAD) remains notably absent. CAD is the industry standard for 3D modeling and plays a vital role in the design and development of products across different industries. As the complexity of modern designs increases, the potential for LLMs to enhance and streamline CAD workflows presents an exciting frontier. This article presents the first systematic survey exploring the intersection of LLMs and CAD. We begin by outlining the industrial significance of CAD, highlighting the need for AI-driven innovation. Next, we provide a detailed overview of the foundation of LLMs. We also examine both closed-source LLMs as well as publicly available models. The core of this review focuses on the various applications of LLMs in CAD, providing a taxonomy of six key areas where these models are making considerable impact. Finally, we propose several promising future directions for further advancements, which offer vast opportunities for innovation and are poised to shape the future of CAD technology. Github: https://github.com/lichengzhanguom/LLMs-CAD-Survey-Taxonomy
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: CSURjournalvolume: 37journalnumber: 4article: 111publicationmonth: 5ccs: Applied computing Computer-aided designccs: Information systems Language modelsccs: Computing methodologies Natural language processingccs: Computing methodologies Computer vision tasks
@@ 1. Introduction
Figure 1. The taxonomy of this review.
[…]
Taxonomy of this review.
Language is a fundamental aspect of human intelligence, as essential as vision. For decades, researchers have strived to endow machines with the ability to reason and communicate in natural language (Zhao et al., 2023). In recent years, this vision has come closer to reality with the advent of Large Language Models (LLMs), which have demonstrated remarkable progress across a wide range of tasks (Shen et al., 2023). A key insight driving LLMs is that knowledge about the world can be effectively captured and represented through large-scale language modeling. This enables the construction of general-purpose models that can tackle a variety of problems via pre-training and task-specific fine-tuning, alignment, or prompting—bypassing the need for extensive domain-specific training from scratch (Xiao and Zhu, 2025).
@@
LLMs are typically built upon the Transformer architecture (Vaswani et al., 2017), with a vastly larger number of parameters compared to earlier Pre-trained Language Models (PLMs). This increase in scale has resulted in significantly improved capabilities in language understanding and generation. The release of ChatGPT (Schulman et al., 2022) marked a major milestone, showcasing unprecedented performance. Since then, many powerful LLMs have been introduced, including GPT-4 (Achiam et al., 2023) and GPT-4V (OpenAI, 2023) by OpenAI; LLaMA (Touvron et al., 2023a), LLaMA-2 (Touvron et al., 2023b), and LLaMA-3 (Grattafiori et al., 2024) by Meta; and Gemini (Team et al., 2023), PaLM (Chowdhery et al., 2023), PaLM-2 (Anil et al., 2023), and GLaM (Du et al., 2022) by Google, among others. Parallel to these advancements in LLMs, Computer-Aided Design (CAD) has remained a cornerstone technology in engineering and industrial design. CAD refers to the use of computers to assist in the creation, modification, analysis, or optimization of a design (Sarcar et al., 2008; Meinders et al., 2008; Gujarathi and Ma, 2011; Quintana et al., 2010). It is widely used to produce 2D drawings and 3D models of physical products, supporting applications in fields such as architecture, automotive design, manufacturing, and 3D printing. CAD enhances design precision, facilitates rapid iteration, and reduces both development time and cost. Moreover, CAD systems preserve essential design information—such as geometry, dimensions, and structural details—in standardized file formats, promoting reusability and collaboration.
Despite the growing popularity of LLMs and the critical importance of CAD in modern industry, no survey article to date has systematically explored the intersection of these two domains. While reviewing efforts have been made for various LLM applications (Zhao et al., 2023; Hadi et al., 2023; Minaee et al., 2024; Naveed et al., 2023; Yin et al., 2024; Kumar, 2024; Annepaka and Pakray, 2024; Hadi et al., 2024; Shen et al., 2023; Hou et al., 2024), a comprehensive survey of exploiting LLMs for CAD still remains missing.
@@
LLMs have already demonstrated transformative potential in various generative and analytical tasks, including image synthesis (Brade et al., 2023; Koh et al., 2023), text generation (Wu, 2024; Li et al., 2024c), and code completion (Cheng et al., 2025; Nejjar et al., 2025). This makes them a promising candidate for advancing CAD through automation, intelligent design assistance, and semantic understanding of design specifications. Hence, a dedicated review of this emerging research area is both timely and essential. It has the potential to offer invaluable insights for researchers and practitioners seeking to integrate LLMs into CAD workflows.
This article presents the first comprehensive survey of research at the intersection of LLMs and CAD. Our objective is to consolidate recent developments, identify trends, and highlight opportunities for future research in this direction. In addition to surveying LLM applications in CAD, we also extend the scope of prior LLM surveys by incorporating the latest generation of models. Specifically, our survey makes the following key contributions.
@@ 2. CAD in Industry
Computer-Aided Design involves the use of computers or workstations to assist in the creation and modification of designs, replacing traditional pencil drawings with precise digital sketches. CAD enhances designers’ productivity, improves design quality, and streamlines communication through clear documentation. Notably, CAD tools can reduce design and prototyping time by 30% to 50%, particularly in iterative design processes (Camburn et al., 2017).
CAD applications span a wide range of sectors, including automotive, aerospace and defense, architecture, manufacturing, electronics, healthcare, consumer goods, energy and utilities, fashion and textiles, entertainment and media, marine and shipbuilding, agriculture and heavy equipment, education and research, jewelry and art [^1] https://www.linkedin.com/pulse/cad-industry-past-present-future-iseeq-ndppf/.
As noted in (Alin and STĂNCIOIU, 2024), CAD is especially important in engineering and manufacturing, buildings and construction, medicine and biotechnology, industrial design and consumer products, and media and entertainment. Similarly, (Ikubanni et al., 2022) identifies its applications in architecture, electronics, shipbuilding, aerospace, textile industry, and education. From these sources, we can observe that CAD is extensively exploited across a variety of practical design related domains.
Taking architectural design as a case in point, CAD is used to model buildings and infrastructure with high precision. Architects and designers can generate detailed 3D models, explore alternative design options, develop construction plans, and collaborate effectively with structural engineers and other project stakeholders (Alin and STĂNCIOIU, 2024). CAD also enables virtual beautification of building designs, helping to reflect realistic expectations visually (Ikubanni et al., 2022).
The global 3D CAD software market reflects this widespread adoption. This market is projected to grow from USD 13.40 billion in 2025 to USD 24.23 billion by 2034222https://www.towardspackaging.com/insights/3d-cad-software-market-sizing?utm_source=chatgpt.com. This growth is largely driven by technological advancements. In 2023, North America led the global CAD software market, while the Asia-Pacific region was forecasted to experience significant growth through 2034. It is noteworthy that the on-premises deployment model accounted for the largest market share in 2023. The large enterprise segment is anticipated to expand significantly between 2025 and 2034. Moreover, the AEC (Architecture, Engineering, and Construction) segment dominated the CAD software market in 2023. These facts present a clear picture of industrial significance of CAD based design.
@@ 3. Foundations of LLMs
In this section, we discuss foundational concepts related to LLMs in an accessible manner. Our aim is to introduce the relevant terms and concepts to non-LLM experts to bridge the gap between LLM and CAD research communities. This section is not aimed at discussing the details of these concepts. Readers interested in technical details are referred to (Naveed et al., 2023; Zhao et al., 2023).
@@ 3.1.1. Tokens and Context Length
In the context of LLMs, tokens are the basic units of text that the model reads, processes, and generates. Tokens are pieces of text, like words, subwords, word pieces, or characters depending on the tokenizer.
The context window, also known as context length, refers to the maximum number of tokens an LLM can see or process at once when generating or understanding text.
@@ 3.1.4. Fine-tuning LLMs
LLMs are first pre-trained on a large corpus of textual data. This pre-training achieves general-purpose language understanding and generation capabilities. Commonly, this is followed by fine-tuning these models to solve specific natural language processing (NLP) tasks. This fine-tuning involves further training of the model on a limited data, which is available for the downstream task. However, fine-tuning only changes the model parameters slightly, thereby making the process of adaptation computationally less expensive. Since the fine-tuning process may also use instruction-following data, such fine-tuning is also known as instruction fine-tuning.
To enable instruction-following, datasets containing various instructions and corresponding responses are required. Scaling the number of such tasks for fine-tuning generally improves model performance (Chung et al., 2024). Unlike pre-training, which may require billions or trillions of samples, fine-tuning can be performed with tens or hundreds of thousands of high-quality samples (Chen et al., 2023). Fine-tuning plays a central role in enabling and enhancing this versatility, and ongoing research continues to improve fine-tuning techniques to make LLMs more efficient and effective.
@@ 3.1.5. Aligning LLMs
Alignment refers to the process of guiding LLMs to behave in accordance with human intentions and ethical standards. This often involves incorporating human feedback, labeled data, or explicitly defined preferences into the training process. Alignment is essential to ensure responsible and safe artificial intelligence (AI) behavior.
Typically, alignment follows two main steps after initial pre-training:
@@
    Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022): In this phase, alignment is treated as a reinforcement learning (RL) problem (Shakya et al., 2023). A reward model—representing the environment—evaluates outputs based on human feedback, while the LLM acts as the agent being optimized to maximize these rewards.
@@ 3.1.6. Prompting LLMs
Prompting plays a key role in effective utilization of LLMs, as it requires no additional training or fine-tuning. LLMs are highly versatile once they are pre-trained on large-scale datasets, and careful prompting can strongly exploit this versatility. Instead of building task-specific systems, users can simply provide well-crafted prompts to perform a wide range of tasks. Consequently, prompt engineering (Perez et al., 2021) has become an active area of research within the NLP community. Owing to the benefits of prompting, LLMs are also known for their zero-shot learning—the ability to perform tasks they were not explicitly trained on.
Another relevant concept in this regard is in-context learning (ICL). In ICL, a well-trained LLM is provided an example of the input-output mapping in the prompt itself during the model inference stage. The model tries to understand the semantics behind this mapping and generalize it to the query in the prompt. The benefit of ICL is that it does not require further training or fine-tuning of the model on the new input-output mapping.
As well-known, LLMs have still been reported to face challenges in tasks that require arithmetic reasoning (Yang et al., 2024a) and commonsense reasoning (Ismayilzada et al., 2023). This shortcoming can be helped with ICL. Nevertheless, the ‘reasoning’ requirements of such tasks call for more sophisticated solutions. Hence, to incorporate reasoning capabilities, researchers use Chain-of-Thought (CoT) prompting, which encourages models to solve problems by breaking them down into a series of intermediate reasoning steps. This mimics human-like cognitive processes and has shown notable improvements, particularly in complex mathematical reasoning tasks (Wang et al., 2024).
There are three common forms of CoT prompting used in ICL:
@@ 3.2.1. Data Source
  • Privacy concerns: When utilizing large-scale datasets, protecting sensitive information is a key concern. LLMs can be fine-tuned to detect and refuse prompts that might lead to data leakage (Wu et al., 2023c).
@@ 3.2.3. Distributed Training
Training LLMs at scale demands significant computational resources, typically provided by distributed systems. To that end, various forms of parallelism are employed, including; data parallelism, model parallelism, tensor parallelism and pipeline parallelism. For instance, LLaMA-3 (405 billion (B) parameters) (Grattafiori et al., 2024) was trained on up to 16,000 H100 GPUs, with each server containing 8 GPUs. Communication across servers was managed using distributed parallelism.
@@ 3.2.4. Scaling Laws
Scaling laws (Kaplan et al., 2020) describe the empirical relationship between LLM performance and key training variables such as model size, dataset size, and computational budget. Research shows that performance consistently improves with increased data—even having the scale of trillions of tokens, and larger models. These scaling laws serve as a blueprint for LLM development, helping researchers make informed decisions about how to allocate resources effectively.
@@ 4.1.1. The GPT Family (OpenAI)
GPT-3 (Brown et al., 2020): Often regarded as the first true LLM, GPT-3 is an autoregressive model with 175 billion parameters. It shares the same architecture as GPT-2 but introduces sparse attention and adopts gradient noise scaling (McCandlish et al., 2018) for better training. GPT-3 is significantly larger than earlier PLMs and exhibits emergent capabilities not seen in smaller models.
CODEX (Chen et al., 2021): A descendant of GPT-3, Codex is fine-tuned on code corpora from GitHub and is optimized for generating code from natural language prompts.
WebGPT (Nakano et al., 2021): Another GPT-3 variant, WebGPT is trained to answer open-ended questions by browsing web pages, enhancing its ability to perform web-based information retrieval.
InstructGPT (Ouyang et al., 2022): This version aligns the model with human intent by fine-tuning it using datasets containing human-written demonstrations and preferences. It shows improved truthfulness and reduced toxicity.
GPT-3.5: Built on GPT-3 and Codex, GPT-3.5 benefits from fine-tuning on code and improved instruction-following as well as RLHF. GPT-3.5 Turbo (Belfathi et al., 2023) is a faster, more cost-effective version optimized for chat.
ChatGPT (Schulman et al., 2022): A major milestone in LLMs, ChatGPT was fine-tuned with human preference data sourced from a wide range of texts (e.g., Wikipedia, books, websites, scientific papers, articles and news media (Welsby and Cheung, 2023)). It excels at delivering engaging, natural conversations (Waghmare, 2023) and completing diverse tasks such as translation, summarization, and Q&A (Shafik, 2024).
@@
GPT-4V (OpenAI, 2023): GPT-4V focuses on safely deploying the visual capabilities of GPT-4, offering strong performance in a variety of vision tasks.
GPT-4o (Hurst et al., 2024): GPT-4o is an “omni” model that handles any combination of text, audio, image, and video as input and output. It brings powerful multimodal capabilities with an emphasis on audio.
OpenAI o1 and o3-mini (Jaech et al., 2024; OpenAI, 2025b):
  • o1 uses a novel optimization technique and a tailored dataset, supporting visual reasoning.
  • o3-mini is a lightweight model optimized for reasoning and cost-efficiency, though it does not support vision.
GPT-4.5 (OpenAI, 2025a): Building on the advancements of GPT-4o, GPT-4.5 offers a more natural interaction experience. It features a broader knowledge base, a stronger alignment with user intent, and improved emotional intelligence, allowing it to better understand and respond to emotional cues. Additionally, GPT-4.5 reduces the occurrence of hallucinations, ensuring more accurate and reliable responses.

### v2

@@ Abstract.
Large Language Models (LLMs) have seen rapid advancements in recent years, with models like ChatGPT and DeepSeek, showcasing their remarkable capabilities across diverse domains. While substantial research has been conducted on LLMs in various fields, a comprehensive review focusing on their integration with Computer-Aided Design (CAD) remains notably absent. CAD is the industry standard for 3D modeling and plays a vital role in the design and development of products across different industries. As the complexity of modern designs increases, the potential for LLMs to enhance and streamline CAD workflows presents an exciting frontier. This article presents the first systematic survey exploring the intersection of LLMs and CAD. We begin by outlining the industrial significance of CAD, highlighting the need for Artificial Intelligence (AI)-driven innovation. Next, we provide a detailed overview of the foundation of LLMs. We also examine both closed-source LLMs as well as publicly available models. The core of this review focuses on the various applications of LLMs in CAD, providing a taxonomy of six key areas where these models are making considerable impact. We also provide a comprehensive study of CAD evaluation, reviewing existing methods and metrics in detail. In our analysis, we also examine common data modalities, model usage trends, dataset sources, and industrial application domains to provide a well-rounded picture of the field. Finally, we propose several promising future directions for further advancements, which offer vast opportunities for innovation and are poised to shape the future of CAD technology. Github: https://github.com/lichengzhanguom/LLMs-CAD-Survey-Taxonomy
@@
copyright: acmlicensedjournalyear: 2026doi: 10.1145/3787499journal: CSURjournalvolume: 58journalnumber: 9article: 225publicationmonth: 2ccs: Applied computing Computer-aided designccs: Information systems Language modelsccs: Computing methodologies Natural language processingccs: Computing methodologies Computer vision tasks
Figure 1. Taxonomy of this review.
@@ 1. Introduction
Language is a fundamental aspect of human intelligence, as essential as vision. For decades, researchers have strived to endow machines with the ability to reason and communicate in natural language (Zhao et al., 2023). In recent years, this vision has come closer to reality with the advent of Large Language Models (LLMs), which have demonstrated remarkable progress across a wide range of tasks (Chang et al., 2024; Zhao et al., 2024; Wan et al., 2024; Huang and Chang, 2023; Wang et al., 2023d; Das et al., 2025; Wu et al., 2024a; Zheng et al., 2025). A key insight driving LLMs is that knowledge about the world can be effectively captured and represented through large-scale language modeling. This enables the construction of general-purpose models that can tackle a variety of problems via pre-training and task-specific fine-tuning, alignment, or prompting, bypassing the need for extensive domain-specific training from scratch (Xiao and Zhu, 2025). LLMs are typically built upon the Transformer architecture (Vaswani et al., 2017), with a vastly larger number of parameters compared to earlier Pre-trained Language Models (PLMs). This increase in scale has resulted in significantly improved capabilities in language understanding and generation. The release of ChatGPT (Schulman et al., 2022) marked a major milestone, showcasing unprecedented performance. Since then, many powerful LLMs have been introduced, including GPT-4 (Achiam et al., 2023) and GPT-4V (OpenAI, 2023b) by OpenAI, LLaMA (Touvron et al., 2023a), LLaMA-2 (Touvron et al., 2023b), and LLaMA-3 (Grattafiori et al., 2024) by Meta, and GLaM (Du et al., 2022), PaLM (Chowdhery et al., 2023), PaLM-2 (Anil et al., 2023), and Gemini (Team et al., 2023) by Google, among others. Parallel to these advancements in LLMs, Computer-Aided Design (CAD) has remained a cornerstone technology in engineering and industrial design. CAD refers to the use of computers to assist in the creation, modification, analysis, or optimization of a design (Sarcar et al., 2008; Meinders et al., 2008; Gujarathi and Ma, 2011; Quintana et al., 2010). It is widely used to produce 2D drawings and 3D models of physical products, supporting applications in fields such as architecture, automotive design, manufacturing, and 3D printing. CAD enhances design precision, facilitates rapid iteration, and reduces both development time and cost. Moreover, CAD systems preserve essential design information such as geometry, dimensions, and structural details in standardized file formats, promoting reusability and collaboration.
@@
Despite the growing popularity of LLMs and the critical importance of CAD in modern industry, no survey article to date has systematically explored the intersection of these two domains. While reviewing efforts have been made for various LLM applications (Hadi et al., 2023; Minaee et al., 2024; Naveed et al., 2025; Yin et al., 2024; Kumar, 2024; Annepaka and Pakray, 2025; Hadi et al., 2024; Shen et al., 2023; Hou et al., 2024), a comprehensive survey of exploiting LLMs for CAD still remains missing. On the other hand, LLMs have already demonstrated transformative potential in various generative and analytical tasks, including image synthesis (Brade et al., 2023; Koh et al., 2023), text generation (Wu, 2024; Li et al., 2024b), and code completion (Cheng et al., 2025; Nejjar et al., 2025). This makes them a promising candidate for advancing CAD through automation, intelligent design assistance, and semantic understanding of design specifications. Hence, a dedicated review of this emerging research area is both timely and essential. It has the potential to offer invaluable insights for researchers and practitioners seeking to integrate LLMs into CAD workflows. This article presents the first comprehensive survey of research at the intersection of LLMs and CAD. Our objective is to consolidate recent developments, identify trends, and highlight opportunities for future research in this direction. In addition to surveying LLM applications in CAD, we also extend the scope of prior LLM surveys by incorporating the latest generation of models. Figure 1 illustrates the structure and organization of this survey. Specifically, our survey makes the following key contributions.
@@
    Study of CAD evaluation: We comprehensively review existing CAD evaluation methods and metrics reported in the literature, integrating diverse approaches to assess CAD models and their practical relevance.
  • 6.
@@ 2. CAD in Industry
CAD involves the use of computers or workstations to assist in the creation and modification of designs, replacing traditional pencil drawings with precise digital sketches. CAD enhances designers’ productivity, improves design quality, and streamlines communication through clear documentation. Notably, CAD tools can reduce design and prototyping time by 30% to 50%, particularly in iterative design processes (Camburn et al., 2017). CAD applications span a wide range of sectors from automotive, aerospace and defense to architecture and manufacturing. Alin et al. (Alin and STĂNCIOIU, 2024) note that CAD is especially important in engineering and manufacturing, buildings and construction, medicine and biotechnology, industrial design and consumer products, and media and entertainment. Similarly, (Ikubanni et al., 2022) identifies CAD applications in electronics, shipbuilding, aerospace, the textile industry, and education. Taken together, these sources illustrate that CAD is extensively employed across a wide variety of practical, design-related domains. Taking architectural design as a case in point, CAD is used to model buildings and infrastructure with high precision. Architects and designers can generate detailed 3D models, explore alternative design options, develop construction plans, and collaborate effectively with structural engineers and other project stakeholders (Alin and STĂNCIOIU, 2024). CAD also enables virtual beautification of building designs, helping to reflect realistic expectations visually (Ikubanni et al., 2022).
The global 3D CAD software market reflects this widespread adoption. This market is projected to grow from USD 13.40 billion in 2025 to USD 24.23 billion by 2034111https://www.towardspackaging.com/insights/3d-cad-software-market-sizing?utm_source=chatgpt.com. This growth is largely driven by technological advancements. In 2023, North America led the global CAD software market, while the Asia-Pacific region was forecasted to experience significant growth through 2034. It is noteworthy that the on-premises deployment model accounted for the largest market share in 2023. The large enterprise segment is anticipated to expand significantly between 2025 and 2034. Moreover, the AEC (Architecture, Engineering, and Construction) segment dominated the CAD software market in 2023. These facts present a clear picture of industrial significance of CAD based design.
@@ 3. Foundations of LLMs
In this section, we discuss foundational concepts related to LLMs in an accessible manner. Our aim is to introduce the relevant terms and concepts to non-LLM experts to bridge the gap between LLM and CAD research communities. This section is not aimed at discussing the details of these concepts. Readers interested in technical details are referred to (Naveed et al., 2025; Zhao et al., 2023).
@@ 3.1.1. Tokens and Context Length
In the context of LLMs, tokens are the basic units of text that the model reads, processes, and generates. Tokens are pieces of text, like words, subwords, word pieces, or characters depending on the tokenizer. The context window, also known as context length, refers to the maximum number of tokens an LLM can see or process at once when generating or understanding text.
@@ 3.1.4. Fine-tuning LLMs
LLMs are first pre-trained on a large corpus of textual data. This pre-training achieves general-purpose language understanding and generation capabilities. Commonly, this is followed by fine-tuning these models to solve specific natural language processing (NLP) tasks. This fine-tuning involves further training of the model on a limited data, which is available for the downstream task. However, fine-tuning only changes the model parameters slightly, thereby making the process of adaptation computationally less expensive. Since the fine-tuning process may also use instruction-following data, such fine-tuning is also known as instruction fine-tuning. To enable instruction-following, datasets containing various instructions and corresponding responses are required. Scaling the number of such tasks for fine-tuning generally improves model performance (Chung et al., 2024). Unlike pre-training, which may require billions or trillions of samples, fine-tuning can be performed with tens or hundreds of thousands of high-quality samples (Chen et al., 2023). Fine-tuning plays a central role in enabling and enhancing this versatility, and ongoing research continues to improve fine-tuning techniques to make LLMs more efficient and effective.
@@ 3.1.5. Aligning LLMs
Alignment refers to the process of guiding LLMs to behave in accordance with human intentions and ethical standards. This often involves incorporating human feedback, labeled data, or explicitly defined preferences into the training process. Alignment is essential to ensure responsible and safe artificial intelligence (AI) behavior. Typically, alignment follows two main steps after initial pre-training:
@@
    Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022): In this phase, alignment is treated as a reinforcement learning (RL) problem (Shakya et al., 2023). A reward model representing the environment evaluates outputs based on human feedback, while the LLM acts as the agent being optimized to maximize these rewards.
@@ 3.1.6. Prompting LLMs
Prompting plays a key role in effective utilization of LLMs, as it requires no additional training or fine-tuning. LLMs are highly versatile once they are pre-trained on large-scale datasets, and careful prompting can strongly exploit this versatility. Instead of building task-specific systems, users can simply provide well-crafted prompts to perform a wide range of tasks. Consequently, prompt engineering (Perez et al., 2021) has become an active area of research within the NLP community. Owing to the benefits of prompting, LLMs are also known for their zero-shot learning ability to perform tasks they were not explicitly trained on. Another relevant concept in this regard is in-context learning (ICL). In ICL, a well-trained LLM is provided an example of the input-output mapping in the prompt itself during the model inference stage. The model tries to understand the semantics behind this mapping and generalize it to the query in the prompt. The benefit of ICL is that it does not require further training or fine-tuning of the model on the new input-output mapping. As well-known, LLMs have still been reported to face challenges in tasks that require arithmetic reasoning (Yang et al., 2024b) and commonsense reasoning (Ismayilzada et al., 2023). This shortcoming can be helped with ICL. Nevertheless, the ‘reasoning’ requirements of such tasks call for more sophisticated solutions. Hence, to incorporate reasoning capabilities, researchers use Chain-of-Thought (CoT) prompting, which encourages models to solve problems by breaking them down into a series of intermediate reasoning steps. This mimics human-like cognitive processes and has shown notable improvements, particularly in complex mathematical reasoning tasks (Wang et al., 2024a). There are three common forms of CoT prompting used in ICL:
@@ 3.2.1. Data Source
  • Privacy and Copyright Concerns: When utilizing large-scale datasets, protecting sensitive information and respecting copyright are key considerations. LLMs can be fine-tuned to detect and refuse prompts that might lead to data leakage (Wu et al., 2023c) or copyright violations (Karamolegkou et al., 2023).
@@ 3.2.3. Distributed Training
Training LLMs at scale demands significant computational resources, typically provided by distributed systems. To that end, various forms of parallelism are employed, including data parallelism, model parallelism, tensor parallelism and pipeline parallelism. For instance, the 405 billion (B) parameter version of LLaMA-3 (Grattafiori et al., 2024) was trained on up to 16000 H100 GPUs, with each server containing 8 GPUs. Communication across servers was managed using distributed parallelism.
@@ 3.2.4. Scaling Laws
Scaling laws (Kaplan et al., 2020) describe the empirical relationship between LLM performance and key training variables such as model size, dataset size, and computational budget. Research shows that performance consistently improves with increased data, even at the scale of trillions of tokens, and larger models. These scaling laws serve as a blueprint for LLM development, helping researchers make informed decisions about how to allocate resources effectively.